what is 2y=6x+4 and 2x+y=7 using substitution ?
Answer by macston(5194)

2y=6x+4 divide both sides by 2
y=3x+2 Use this to substitute for y in second equation
2x+y=7
2x+(3x+2)=7
5x+2=7 Subtract 2 from each side
5x=5
x=1 ANSWER X is 1
2y=6x+4
2y=6+4
2y=10
y=5 ANSWER Y is 5
CHECK
2x+y=7
2(1)+5=7
2+5=7
7=7